# Lectures on Gauge theories and Many-Body systems

### v1

@@ 7.7 Codimension 2 vortex-type operators in four dimensional Yang-Mills theory
For the algebraic geometric approach relating parabolic bundles to orbifolds we use below, see [biswas].
@@ 7.9 $\hat{A}_{0}$-$q$-character
is holomorphic in the $x$-variable. This observable is called a fundamental $qq$-character for $\hat{A}_{0}$-theory (or order operator) ([Nekrasov2015BPSCFTCN]).
@@ 7.10 The order operator as pushforward
The formula (7.1) has the interpretation in terms of a slightly more general statistical model, a particular instance of gauge origami [Nekrasov:2016ydq]:
The measured space is the set of pairs of multipartitions $\Pi^{N}\times\Pi^{M}$ equipped with the product measure $\mu_{(\lambda^{(\alpha)})\times(\lambda^{(\beta)})}=\mu^{1}_{(\lambda^{(\alpha)})}\times\mu^{2}_{(\lambda^{(\beta)})}$. Each factor is the measure introduced earlier:
@@
([Nekrasov2015BPSCFTCN]) Define the observable
@@ 7.11 Disorder operators: Surface defects
Analogously to Wilson loop operators in Chern–Simons theory one can explore codimension-two defects in 4d Yang–Mills ([Nekrasov:2021tik]). This leads to the study of sheaves equipped with parabolic structure. For example, for a sheaf $\mathcal{E}$ on $\mathbb{C}^{2}$ a parabolic structure along the $D=\{z_{2}=0\}\subset\mathbb{C}^{2}$-plane is the filtration with subsheaves
@@
$\Omega$ $\mapsto e^{\frac{2\pi\sqrt{-1}\omega}{N}}.$
It allows us to make no distinction between $\hat{\Gamma}$ and $\{0,..,N-1\}$ in the following. Let us fix an arbitrary bijection $c:\{1,...,N\}\rightarrow\{0,...,N-1\}$. For each multipartition $(\lambda^{(\alpha)})$ we dye its every Young diagram $\lambda^{(\alpha)}$ in colors of $\hat{\Gamma}$. Boxes in the $j$’s column have by definition the same color $c(\alpha)+j-1\mod N\in\mathbb{Z}/N\mathbb{Z}$. For every $\omega\in\{0,...,N-1\}$ we define an observable $k_{\omega}=k_{\omega}[(\lambda^{(\alpha)})]$ as the total length
@@
and a truncation operator $f\mapsto f^{\delta}$, acting on functions $f:L\rightarrow\mathbb{C}$:
$f^{\delta}(x)=\begin{cases}f(x),[x]=0\\ 1,\;\text{otherwise}.\end{cases}$ (7.50)
@@
$\mu^{\rm orb}_{(\lambda^{(\alpha)})}=\prod\limits_{\omega=0}^{N-1}\mathfrak{q}_{\omega}^{k_{\omega}}\prod\limits_{\alpha,\beta=1}^{N}$ $\prod\limits_{(i,j)\in\lambda^{(\beta)}}\frac{\theta^{\delta}(\varepsilon_{3}+\varepsilon_{1}(i-\lambda_{j}^{(\beta)t})+\varepsilon_{2}(1+\lambda_{i}^{(\alpha)}-j)+a_{\alpha}-a_{\beta})}{\theta^{\delta}(\varepsilon_{1}(i-\lambda_{j}^{(\beta)t})+\varepsilon_{2}(1+\lambda_{i}^{(\alpha)}-j)+a_{\alpha}-a_{\beta})}\cdot$
$\prod\limits_{(i,j)\in\lambda^{(\alpha)}}\frac{\theta^{\delta}(\varepsilon_{3}+\varepsilon_{1}(\lambda_{j}^{(\alpha)t}+1-i)+\varepsilon_{2}(j-\lambda_{i}^{(\beta)})+a_{\alpha}-a_{\beta})}{\theta^{\delta}(\varepsilon_{1}(\lambda_{j}^{(\alpha)t}+1-i)+\varepsilon_{2}(j-\lambda_{i}^{(\beta)})+a_{\alpha}-a_{\beta})}$
@@
$\pi:\Pi^{N}$ $\rightarrow\Pi^{N}$
$(\lambda^{(\alpha)})$ $\mapsto(\Lambda^{(\alpha)}),\;\;\Lambda_{j}^{(\alpha)t}=\lambda_{\alpha+N(j-1)}^{(\alpha)t}.$
@@
$\pi_{*}\mu^{\rm orb}$ $=\mu(\vec{a},\tau,{{\varepsilon}}_{3};\varepsilon_{1},\varepsilon_{2})\times\mathcal{S}(\vec{w}),$
$e^{2\pi\sqrt{-1}\tau}$ $=\mathfrak{q}=\prod\limits_{\omega\in\hat{\Gamma}}\mathfrak{q}_{\omega},\;\;\vec{w}=[w_{0}:...:w_{N}]$
$\frac{w_{1}}{w_{0}}$ $=\mathfrak{q}_{0},\penalty 10000\ ...\penalty 10000\ ,\frac{w_{N-1}}{w_{N-2}}=\mathfrak{q}_{N-2},\frac{w_{N}}{w_{N-1}}=\mathfrak{q}^{-1}\mathfrak{q}_{N-1}.$
@@ 8 Indirect gauge theory — integrable systems correspondence
In this section we demonstrate a relation ([Fock:1999ae], [Grekov:2023fek], [Grekov:2024fyh]) between statistical models originating from gauge theories and integrable systems on a particular example. Among many applications of this relation is the classical/quantum correspondence including Bethe ansatz and its connection to geometric Langlands program first observed in [Feigin:1994in]. We refer the reader to [Etingof:2023drx, Jeong:2023qdr, Jeong:2024hwf] for modern developments and their place in the BPS/CFT correspondence.
We consider (some specification of) $\hat{A}_{r}$ theory which is a generalization of previously considered $\hat{A}_{0}$ - theory.
@@ 8.1 $\hat{A}_{r+1}$-theory with surface defect
$\Pi^{N_{0}}\times...\times\Pi^{N_{r+1}}=\{(\lambda^{(\alpha,s)})_{\alpha=1,...,N_{s}}^{s=0,...,r+1}:\lambda^{(\alpha,s)}\in\Pi\}.$
@@
$(a_{\alpha,s})_{\alpha=1,...,N_{s}}^{s=0,...,r+1},\in\mathbb{C}^{N_{0}}\times...\times\mathbb{C}^{N_{r+1}},$
@@
where $\mathfrak{X}_{\lambda}(\omega,s,x,(\lambda^{(\alpha,s)}),\varepsilon_{1},...,\varepsilon_{4})$ are some explicit (but slightly complicated) expressions of parameters. In the following we are interesting in $A_{r}$-theory, which might be considered as a degeneration of $\hat{A}_{r}$-theory.
@@ 8.4 Lax operator
This is the Lax operator of Garnier–Gaudin system. The matrix valued $1$-form $\phi(z)=\hat{L}(z)\frac{dz}{z}$ is the Higgs field for Hitchin system on $\mathbb{CP}^{1}$ with $r+3$ punctures with general residues at 0 and $\infty$ and rank $1$ residues at $z_{i}$’s. Note that there is a similar construction for $\hat{A}_{0}$-theory involving infinite products ([Grekov:2023fek], [Grekov:2024fyh]). The corresponding $\hat{L}(z)$ is the Lax operator of elliptic Calogero–Moser system.
@@ 9.1 Problem 1
$L^{(1)}_{ij}=p_{i}\delta_{ij},\;\;\;L^{(2)}_{ij}=(1-\delta_{ij})\frac{\sqrt{-1}\nu}{x_{i}-x_{j}},$
$A^{(1)}_{ij}=(1-\delta_{ij})\frac{i\sqrt{-1}\nu}{(x_{i}-x_{j})^{2}},\;\;\;A^{(2)}_{ij}=-\delta_{ij}\sqrt{-1}\nu\sum\limits_{k\neq i}\frac{1}{(x_{i}-x_{k})^{2}}.$
@@
$[A^{(1)},L^{(1)}]_{ij}=(1-\delta_{ij})\frac{\sqrt{-1}\nu(p_{j}-p_{i})}{(x_{i}-x_{j})^{2}},$
@@
$[A,L]_{ij}=[A^{(1)},L^{(1)}]_{ij}=(1-\delta_{ij})\frac{\sqrt{-1}\nu(p_{j}-p_{i})}{(x_{i}-x_{j})^{2}}+2\nu^{2}\delta_{ij}\sum\limits_{k\neq i}\frac{1}{(x_{i}-x_{k})^{3}}.$ (9.1)
@@
$\dot{L}_{ij}=\dot{p}_{i}\delta_{ij}+(1-\delta_{ij})\sqrt{-1}\nu\frac{\dot{x}_{j}-\dot{x}_{i}}{(x_{i}-x_{j})^{2}}.$ (9.2)
@@ 9.2 Problem 2
$(x_{j}-x_{i})P_{ij}+\sqrt{-1}z_{i}z_{j}-\sqrt{-1}\nu\delta_{ij}=0.$
@@ 9.3 Problem 3
$\begin{cases}\partial_{\bar{z}}E_{ij}-(\alpha_{i}-\alpha_{j})E_{ij}+(\delta_{ij}\nu-v_{i}w_{j})2\pi\sqrt{-1}\delta^{(2)}(z,\bar{z})=0\\ E_{ij}(z+1)=E_{ij}(z+\tau)=E_{ij}(z)\end{cases}$ (9.3)
@@
In fact $L_{ij}$ is uniquely defined by these constraints and so we come up to a Krichever’s Lax matrix for the Elliptic Calogero-Moser system ([Kr])
@@ 9.9 Problem 9
Consider the operator $\hat{D}_{0}=\prod\limits_{i=0}^{\overset{\longrightarrow}{r}}\big(1-\hat{C}_{z}\hat{Z}_{i}\big)$. It degenerates when $\det(1-z^{-1}\hat{C}_{z}\hat{Z}_{i})=0$ for some $i$. By the previous problem it happens when $z=\det\hat{Z}_{i}=z_{i}$. Since $z_{i}\neq z_{j}$ for different $i$ and $j$ the operators $1-\hat{C}_{z_{i}}\hat{Z}_{j}$ are invertible and thus the dimension of $\ker\hat{D}_{0}(z_{0})$ is equal to the dimension of $\ker(1-\hat{C}_{z_{i}}\hat{Z}_{i})$. From the explicit form (9.8) we see that the first N1N-1 rows of $1-\hat{C}_{z}\hat{Z}_{i}$ are always independent so $\mathrm{rk}(1-\hat{C}_{z_{0}}\hat{Z}_{i})=N-1$ and hence $\dim\ker(1-\hat{C}_{z}\hat{Z}_{i})=1$. Now consider the product
@@
$\mathrm{im}\mathrm{Res}_{z=z_{0}}\hat{L}(z)\subset\ker\hat{D}_{0}(z_{0}).$

### v2

@@ 7.7 Codimension 2 vortex-type operators in four dimensional Yang–Mills theory
For the algebraic geometric approach relating parabolic bundles to orbifolds we use below, see [Bis97].
@@ 7.9 $\widehat{A_{0}}-$$q$-character
is holomorphic in the $x$-variable. This observable is called a fundamental $qq$-character for $\hat{A}_{0}$-theory (or order operator) ([Nek15]).
@@ 7.10 The order operator as pushforward
The formula (7.1) has the interpretation in terms of a slightly more general statistical model, a particular instance of gauge origami [Nek18]:
The measured space is the set of pairs of multi–partitions $\Pi^{N}\times\Pi^{M}$ equipped with the product measure $\mu_{(\lambda^{(\alpha)})\times(\lambda^{(\beta)})}=\mu^{1}_{(\lambda^{(\alpha)})}\times\mu^{2}_{(\lambda^{(\beta)})}$. Each factor is the measure introduced earlier:
@@
([Nek15]) Define the observable
@@ 7.11 Disorder operators: Surface defects
Analogously to Wilson loop operators in Chern–Simons theory one can explore codimension-two defects in 4d Yang–Mills ([NT22]). This leads to the study of sheaves equipped with parabolic structure. For example, for a sheaf $\mathcal{E}$ on $\mathbb{C}^{2}$ a parabolic structure along the $D=\{z_{2}=0\}\subset\mathbb{C}^{2}$-plane is the filtration with subsheaves
@@
$\Omega$ $\mapsto e^{\frac{2\pi{\mathrm{i}}\omega}{N}}.$
It allows us to make no distinction between $\hat{\Gamma}$ and $\{0,..,N-1\}$ in the following. Let us fix an arbitrary bijection $c:\{1,...,N\}\rightarrow\{0,...,N-1\}$. For each multi–partition $(\lambda^{(\alpha)})$ we dye its every Young diagram $\lambda^{(\alpha)}$ in colors of $\hat{\Gamma}$. Boxes in the $j$’s column have by definition the same color $c(\alpha)+j-1\mod N\in\mathbb{Z}/N\mathbb{Z}$. For every $\omega\in\{0,...,N-1\}$ we define an observable $k_{\omega}=k_{\omega}[(\lambda^{(\alpha)})]$ as the total length
@@
and a projection operator $f\mapsto f^{\delta}$, acting on functions $f:L\rightarrow\mathbb{C}$:
$f^{\delta}(x)=\begin{cases}f(x),\ &[x]=0\\ \ 1,\ &\text{otherwise}.\end{cases}$ (7.50)
@@
$\mu^{\rm orb}_{(\lambda^{(\alpha)})}=\prod\limits_{\omega=0}^{N-1}\mathfrak{q}_{\omega}^{k_{\omega}}\prod\limits_{\alpha,\beta=1}^{N}$ $\prod\limits_{(i,j)\in\lambda^{(\beta)}}\frac{\theta^{\delta}\left(\varepsilon_{3}+\varepsilon_{1}(i-\lambda_{j}^{(\beta)t})+\varepsilon_{2}\left(1+\lambda_{i}^{(\alpha)}-j\right)+a_{\alpha}-a_{\beta}\right)}{\theta^{\delta}\left(\varepsilon_{1}\left(i-\lambda_{j}^{(\beta)t}\right)+\varepsilon_{2}\left(1+\lambda_{i}^{(\alpha)}-j\right)+a_{\alpha}-a_{\beta}\right)}\ \times$
$\prod\limits_{(i,j)\in\lambda^{(\alpha)}}\frac{\theta^{\delta}\left(\varepsilon_{3}+\varepsilon_{1}(\lambda_{j}^{(\alpha)t}+1-i)+\varepsilon_{2}\left(j-\lambda_{i}^{(\beta)}\right)+a_{\alpha}-a_{\beta}\right)}{\theta^{\delta}\left(\varepsilon_{1}\left(\lambda_{j}^{(\alpha)t}+1-i\right)+\varepsilon_{2}\left(j-\lambda_{i}^{(\beta)}\right)+a_{\alpha}-a_{\beta}\right)}$
@@
$\pi\,:\,\Pi^{N}\ \longrightarrow\Pi^{N}\,$,
${\pi}\,:\,(\lambda^{(\alpha)})$ $\mapsto(\Lambda^{(\alpha)})\,,\;\;\Lambda_{j}^{(\alpha)t}=\lambda_{\alpha+N(j-1)}^{(\alpha)t}.$
@@
$\pi_{*}\mu^{\rm orb}$ $=\mu({\bf a},\tau,{{\varepsilon}}_{3};{\varepsilon}_{1},{\varepsilon}_{2})\times\mathcal{S}(\vec{w}),$
$e^{2\pi{\mathrm{i}}\tau}$ $=\mathfrak{q}=\prod\limits_{\omega\in\hat{\Gamma}}\mathfrak{q}_{\omega}\,,\ \vec{w}=[w_{0}:...:w_{N}]$
$\frac{w_{1}}{w_{0}}$ $=\mathfrak{q}_{0}\,,\,\ldots\,,\,\frac{w_{N-1}}{w_{N-2}}=\mathfrak{q}_{N-2}\,,\,\frac{w_{N}}{w_{N-1}}=\mathfrak{q}^{-1}\mathfrak{q}_{N-1}.$
@@ 8 Indirect gauge theory — integrable systems correspondence
In this section we illustrate, through a particular example, the relation between the statistical models arising from gauge theories and integrable systems. Among the many applications of this relation are the classical/quantum correspondence, including the Bethe ansatz, and the connection to the geometric Langlands program, first observed in [FFR94]. We refer the reader to [EFK24, JLN23, JLN24] for modern developments and their place in the BPS/CFT correspondence.
We consider a particular version of $\hat{A}_{r}$-theory, which generalizes the previously discussed $\hat{A}_{0}$-theory.
@@ 8.1 $\hat{A}_{r+1}$-theory with surface defect
$\Pi^{N_{0}}\times\ldots\times\Pi^{N_{r+1}}=\left\{\,\left(\lambda^{(\alpha,s)}\right)_{\alpha=1,...,N_{s}}^{s=0,...,r+1}\,\Biggr|\ {\lambda}^{(\alpha,s)}\in\Pi\,\right\}.$
@@
$(a_{\alpha,s})_{\alpha=1,...,N_{s}}^{s=0,...,r+1}\in\mathbb{C}^{N_{0}}\times...\times\mathbb{C}^{N_{r+1}},$
@@
where $\mathfrak{X}_{\lambda}(\omega,s,x,(\lambda^{(\alpha,s)}),\varepsilon_{1},...,\varepsilon_{4})$ are some explicit (but slightly complicated) expressions of parameters. In the following we focus on the $A_{r}$-theory, which may be viewed as a degeneration of $\hat{A}_{r}$-theory.
@@ 8.4 Lax operator
This is the Lax operator of Garnier–Gaudin system. The matrix valued $1$-form $\phi(z)=\hat{L}(z)\frac{dz}{z}$ is the Higgs field for Hitchin system on $\mathbb{CP}^{1}$ with $r+3$ punctures with general residues at 0 and $\infty$ and rank $1$ residues at $z_{i}$’s. Note that there is a similar construction for $\hat{A}_{0}$-theory involving infinite products ([GN23], [Gre24]). The corresponding $\hat{L}(z)$ is the Lax operator of elliptic Calogero–Moser system.
@@ 9.1 Problem 1
$L^{(1)}_{ij}=p_{i}\delta_{ij},\;\;\;L^{(2)}_{ij}=(1-\delta_{ij})\frac{{\mathrm{i}}\nu}{x_{i}-x_{j}},$
$A^{(1)}_{ij}=(1-\delta_{ij})\frac{{\mathrm{i}}\nu}{(x_{i}-x_{j})^{2}},\;\;\;A^{(2)}_{ij}=-\delta_{ij}\sum\limits_{k\neq i}\frac{{\mathrm{i}}\nu}{(x_{i}-x_{k})^{2}}.$
@@
$[A^{(1)},L^{(1)}]_{ij}=(1-\delta_{ij})\frac{{\mathrm{i}}\nu(p_{j}-p_{i})}{(x_{i}-x_{j})^{2}},$
@@
$[A,L]_{ij}=[A^{(1)},L^{(1)}]_{ij}=(1-\delta_{ij})\frac{{\mathrm{i}}\nu(p_{j}-p_{i})}{(x_{i}-x_{j})^{2}}+2\nu^{2}\delta_{ij}\sum\limits_{k\neq i}\frac{1}{(x_{i}-x_{k})^{3}}.$ (9.1)
@@
$\dot{L}_{ij}=\dot{p}_{i}\delta_{ij}+(1-\delta_{ij}){\mathrm{i}}\nu\frac{\dot{x}_{j}-\dot{x}_{i}}{(x_{i}-x_{j})^{2}}.$ (9.2)
@@ 9.2 Problem 2
$(x_{j}-x_{i})P_{ij}+{\mathrm{i}}z_{i}z_{j}-{\mathrm{i}}\nu\delta_{ij}=0.$
@@ 9.3 Problem 3
$\begin{cases}\partial_{\bar{z}}E_{ij}-(\alpha_{i}-\alpha_{j})E_{ij}+(\delta_{ij}\nu-v_{i}w_{j})2\pi{\mathrm{i}}\delta^{(2)}(z,\bar{z})=0\\ E_{ij}(z+1)=E_{ij}(z+\tau)=E_{ij}(z)\end{cases}$ (9.3)
@@
In fact $L_{ij}$ is uniquely defined by these constraints and so we come up to a Krichever’s Lax matrix for the Elliptic Calogero–Moser system [KLWZ97]
@@ 9.9 Problem 9
Consider the operator $\hat{D}_{0}=\prod\limits_{i=0}^{\overset{\longrightarrow}{r}}\big(1-\hat{C}_{z}\hat{Z}_{i}\big)$. It degenerates when $\det(1-z^{-1}\hat{C}_{z}\hat{Z}_{i})=0$ for some $i$. By the previous problem it happens when $z=\det\hat{Z}_{i}=z_{i}$. Since $z_{i}\neq z_{j}$ for different $i$ and $j$ the operators $1-\hat{C}_{z_{i}}\hat{Z}_{j}$ are invertible and thus the dimension of ${\mathrm{Ker}}\hat{D}_{0}(z_{0})$ is equal to the dimension of ${\mathrm{Ker}}(1-\hat{C}_{z_{i}}\hat{Z}_{i})$. From the explicit form (9.8) we see that the first N1N-1 rows of $1-\hat{C}_{z}\hat{Z}_{i}$ are always independent so $\mathrm{rk}\,(1-\hat{C}_{z_{0}}\hat{Z}_{i})=N-1$ and hence $\dim{\mathrm{Ker}}(1-\hat{C}_{z}\hat{Z}_{i})=1$. Now consider the product
@@
$\mathrm{im}\,\mathrm{Res}_{z=z_{0}}\,\hat{L}(z)\subset{\mathrm{Ker}}\hat{D}_{0}(z_{0}).$